OpenStudy (anonymous):

If u = (1 + i ) and v = (1 − i ), what is uv? A) 1 B) -4 C) -4i D) 4

OpenStudy (mathmate):

Not sure if your choices are correct, or perhaps there is a typo in the question or answers.

OpenStudy (mathmate):

\((1+i)(1-i)=1^2-i^2=1-(-1)=2\)

OpenStudy (nurali):

u = (1+i ), v=(1−i ) uv=(1+i)(1-i) uv=(1-i+i-i^2) uv=(1-i^2) i^2=-1 uv=(1-(-1)) uv=2